OpenStudy (derpisthewordorisit):

Replace ? with =, >, or < to make the statement true. 4 ∙ 16 – 16 ? 4 ∙ [24 – 2 ∙ (4 + 8)] A. < B. = C. >

Mehek (mehek14):

this time you try for 4 ∙ 16 – 16 do 4 ∙ 16 and then subtract 16

OpenStudy (derpisthewordorisit):

So C? Kinda forgot my least and greatest signs >.>

Mehek (mehek14):

lets see 4*16=64 64-16= 48

OpenStudy (derpisthewordorisit):

Yes, I know I got it. But I meant the least sign?

OpenStudy (derpisthewordorisit):

And the greatest.

Mehek (mehek14):

4 ∙ [24 – 2 ∙ (4 + 8)] 4+8=12 12*2= 24 24-24=0 so its greater than >

OpenStudy (derpisthewordorisit):

OKzy, thanks.

Mehek (mehek14):

yw ^_^

OpenStudy (derpisthewordorisit):

Can you help me with anotah 1?

Mehek (mehek14):

sure